OpenStudy (anonymous):

write an equation in standard form with an x-intercept of 5 and a y-intercept of -4

OpenStudy (anonymous):

X intercept (5,0); y intercept (0,-4) the slope=-4/-5=4/5 y=(4/5)x-4 (4/5)x-y=4